Planw A is traveling at 300 mph ... Plane B at 400 mph. The current time is 1301 and they are 61 miles apart. What time will the two aircraft meet?
Answer by rwm(914)

it depends on where they are going now doesn't it?
assuming they are heading towards each other
they are traveling at a combined speed of 700 mph
700*x=61
x=61/700=.087 hours which is about 5.23 minutes
so they meet at 1306.23